OpenStudy (anonymous):

find the value of sin^-1,0

ganeshie8 (ganeshie8):

say, \(x = \sin^{-1}0 \) take \(\sin\) both sides \(\sin x = \sin~ \sin^{-1} 0\) \(\sin x = 0\) wat value of x makes sinx = 0 ?

OpenStudy (anonymous):

is it 0?

ganeshie8 (ganeshie8):

thats it, sin 0 = 0 also, sin (pi) = 0

ganeshie8 (ganeshie8):

so the answer is 0, pi within the interval [0, 2pi)